OpenStudy (anonymous):

How do you factor completely: 12x^-6y^-3z^4 + 8x^-3y^-2z^-1 - 6x^3y^2z^-3 +10

OpenStudy (anonymous):

incorrect expression ..

OpenStudy (anonymous):

What do you mean it's incorrect?

OpenStudy (turingtest):

8x^-3y^-2z^-1 should be written as (8x^-3)(y^-2)(z^-1) or better yet try the equation feature:\[8x^{-3}y^{-2}z^{-1}\]but even with that I don't see how to factor \[12x^{-6}y^{-3}z^4 + 8x^{-3}y^{-2}z^{-1} - 6x^3y^2z^{-3} +10\] try Wolfram...

OpenStudy (anonymous):

What or who is Wolfram?

OpenStudy (turingtest):

Wolfram Alpha is a site that will factor and integrate things for you. Try to enter your expression (you will need to add parentheses on the exponents first) into http://www.wolframalpha.com/

OpenStudy (turingtest):

and write the word "factor" before it.

OpenStudy (anonymous):

Something like this \[ 2x^{-6}y^{-3}z^4 + 8x^{-3}y^{-2}z^{-1} - 6x^3y^2z^{-3} +10 = -\frac{2 \left(3 x^9 y^5-4 x^3 y z^2-5 x^6 y^3 z^3-z^7\right)}{x^6 y^3 z^3} \]

OpenStudy (anonymous):

So basically the only common factor you can pull out is the "2"?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Is it necessary to find a common denominator for a factoring problem?

OpenStudy (turingtest):

That and you need to recognize that the denominator will be composed if the negative exponents LCM. http://www.wolframalpha.com/input/?i=factor+12x%5E%28-6%29y%5E%28-3%29z%5E4+%2B+8x%5E%28-3%29y%5E%28-2%29z%5E%28-1

OpenStudy (anonymous):

No,but it's a good practice

OpenStudy (turingtest):

composed OF^

OpenStudy (anonymous):

Thanks guys